OpenStudy (albert0898):

A fan is running at 10 revolutions per second. After it is turned off, it's speed decreases at a rate of 75% per second. Determine the number of revolutions completed by the fan after it is turned off.

OpenStudy (amistre64):

what sequence do you think we should apply?

OpenStudy (albert0898):

Infinite Geometric?

OpenStudy (amistre64):

sounds good to me, and how to apply it?

OpenStudy (albert0898):

\[S \infty = \frac{ a }{ 1 - r }\]

OpenStudy (albert0898):

r = .75 a = first term = 10

OpenStudy (albert0898):

I was wondering it that was the correct approach and answer? Or do I use a different formula?

OpenStudy (amistre64):

i dont think r=.75, i could be mistaken tho if it decreases at a rate of k, then it retains 1-k of its speed, right?

OpenStudy (albert0898):

Yes

OpenStudy (albert0898):

So if it decreases .75, then it retains .25

OpenStudy (amistre64):

10 + 10(.25) + 10(.25)(.25) + ...

OpenStudy (amistre64):

so 40/3 sounds fair

OpenStudy (albert0898):

How did you get that though, 40/3?

OpenStudy (amistre64):

10/(1-.25) = 10/(3/4) = 10*4/3

OpenStudy (albert0898):

Ah

OpenStudy (albert0898):

R is the common ratio, so r = .25?

OpenStudy (amistre64):

i would say so. we want to determine how much of the rotations we are retaining. in order to count them.

OpenStudy (albert0898):

This type of problem just happened to be the one I got incorrect on my exam. Thank you for your help!

OpenStudy (amistre64):

accursed exams!! good luck :)

OpenStudy (albert0898):

Thank you! :D